# Emergence of Hidden Capabilities: Exploring Learning Dynamics in Concept Space

### v1

@@ A.1 Synthetic data
Fig. 11 illustrates two datasets with varying concept signals. Fig. 11 (a) illustrates a case where the size concept signal is stronger than the color concept signal, while Fig. 11 (b) illustrates the inverse case where the color concept signal is stronger.
A.2 Model Details
@@
Classifier Free Guidance At training time we drop the conditioning to a $-1.0$ filled null vector $\phi$ with probability 0.2. This choice was made as opposed to the conventional choice of dropping it to the null vector since the null vector was near the interpolation limit of our model in the color subspace of the conditioning. The $w_{cfg}$ parameter is used to estimate the noise in the diffusion process by: $\hat{\epsilon}_{t-1}=f_{\theta}(x_{t}|\phi)+w_{cfg}*(f_{\theta}(x_{t}|h)-f_{\theta}(x_{t}|\phi))$, where $\theta$ denotes parameters of the network $f$.
@@
Diffusion Process Hyperparameters We used the continuous time variational diffusion model [88] framework to (i) keep the sampling step parameter T an inference time hyperparameter and (ii) to allow the model to the adjust its optimal noise schedule for these images, especially since our synthetic images are expected to be different in SNR from natural images. In particular, we initialize our network with $\gamma_{i}=-5.0$ and $\gamma_{f}=10.0$ (see [88] for the definition of $\gamma$) and use a learned linear schedule of $\gamma(t)$. We use a reconstruction loss corresponding to a negative log likelihood from a standard normal with $\sigma=10^{-3}$ centered at the data for the first step of the diffusion process.
@@
Evaluation probe Details. We used a U-Net [93] backbone with 64 output channels followed by a max pooling layer and $n$ 1-layer MLP classifier for each of the $n$ classes to estimate each concept of an image independently. We sample from the same data distribution but with maximal data diversity, i.e., with $s_{i}$ values in App. A.1 maximized within the range allowing perfect classification (no overlap of $z$ between classes). We sample 4096 images per class from the DGP and train the classifier for 10K gradient steps with AdamW [115] and achieve a 100% accuracy on the held out test set. At evaluation phase, we average the classifier softmax output over 5 data generation / model initialization seeds and 32 inference samples to construct the concept space representation of the generations.
Computational Details. We implement our models in PyTorch [116]. A standard model run (e.g., in Sec. 4.2) took $\sim$ 20 minutes on a single NVIDIA A100 40GB GPU. The CelebA runs took $\sim$ 24 hours on the same GPU.
A.3 Training Procedure
@@
The diffusion model was implemented using PyTorch and trained on four Nvidia A100 GPUs. We conducted a hyperparameter search using a validation set, testing batch sizes from 32 to 256, number of channels per layer from 64 to 512, learning rates between $10^{-4}$ and $10^{-3}$, the number of steps in the diffusion process from 100 to 400. We employed the Adam optimizer [117] with $\beta_{1}=0.9$, $\beta_{2}=0.99$, and weight decay of $10^{-5}$.
Appendix B Details on Alternative Protocols for Eliciting Model Capabilities
@@
Our model is trained on a distribution of conditioning centered around a class dependent mean: $p^{i}(z_{j})=\mathcal{U}(m^{i}_{j}-s_{j},m^{i}_{j}+s_{j})$ for each class $i$. We prompt the model with conditioning vectors extrapolated in the direction $\vec{m_{1}}-\vec{m_{0}}$. For instance, assuming the red conditioning was $(0.6,0.4,0.4)$ and the blue conditioning was $(0.4,0.4,0.6)$, we “prompt” the model with $(0.2,0.2,0.8)$. In practice, we use 5 conditionings $(0.4,0.4,0.6)$, $(0.35,0.35,0.65)$, $(0.25,0.25,0.75)$, $(0.15,0.15,0.85)$, $(0.05,0.05,0.95)$, and report the maximum joint accuracy.
@@
Appendix C Additional Results
C.1 Loss and Accuracy versus Concept Space
@@
We plot loss, accuracy, and training trajectory in concept space of the models from Fig. 6. The time at which the model acquires a capability to manipulate the concept color is not evident from the loss or accuracy curves; however, in concept space, it is evident that the concept is learned together with the departure from concept memorization. Benchmarking generative models is a challenging task, still often involving humans in the loop [102, 103]. Our concept space framework suggests that benchmarking out-of-distribution (OOD) generalization can potentially be reduced to monitoring the learning trajectory in concept space.
@@
Figure 12: Loss vs. Accuracy vs. Concept Space. Loss and accuracy do not always tell the full story of what, and even more importantly when, the model is learning a capability. The star symbol in panel (c) shows point of sudden turn and concept learning, i.e., the moment when well defined prompting protocols can elicit the desired output from the model, indicating the model has learned the capability to alter a concept.
C.2 Additional trajectories of learning dynamics (Fig. 13)
@@
Here we show all concept space trajectories for the experiments mentioned in Fig. 4 (a,b), for all classes and color concept signal levels. We find asymmetric behavior for the 01 class and the 10 class when adjusting the color concept signal level. The dynamics of the generations in the training set matches our intuitions. At low color concept signal, we observe that the dynamics fit the size for both 00 and 10, and diverge towards their correct colors.
@@
Figure 13: Concept Space Dynamics for all classes (00, 01, 10, 11). The experiment is identical as in Fig. 4. The [0,1) normalized color concept signal is color coded in every trajectory. Two training data trajectories are shown in gray in the last panel to show concept memorization.
C.3 Experiments with real data: CelebA
@@
To assess our findings on a more realistic dataset, we ran experiments on the CelebA [118] dataset. We selected the Not Male and Smiling features as the two concept dimensions to explore. This choice was motivated by the relatively balanced number of samples in all four classes(00=(Male, Not Smiling), 01=(Male, Smiling), 10=(Not Male, Not Smiling), 11=(Not Male, Smiling)), from which we randomly sampled 30K images in each class. To construct the concept space, we trained a fully convolutional network with a average pooling layer followed by a classification MLP head to classify the two attributes with independent cross entropy losses. (See A.2). We trained the classifier with the AdamW [115] optimizer with learning rate $10^{-3}$ and weight decay $10^{-5}$ for 10K gradient steps. The classifier achieved a final accuracy of respectively $95\%$ and $97\%$ on the held out validation set, which was $10\%$ of the entire dataset.
@@
We trained the same diffusion model (See App. A.2) from our synthetic experiments on 64x64 resized images from the classes (00, 01, 10) and assessed the out-of-distribution (OOD) generalization to 11. We used a color jitter of 0.1 for brightness, contrast and saturation and randomly flipped the images horizontally. As the class attributes are categorical, they were one hot encoded and concatenated to an input conditioning vector of 4 dimensions. The diffusion model was trained for $10^{6}$ gradient steps with a batch size of 64 with the same optimizer as in the main experiments.
@@
The concept space dynamics of generations from the in distribution conditioning and OOD conditioning are shown in Fig. 14. We find similar observations from the concept space trajectories as in our synthetic experiments. Initially, the images corresponding to the class 11 follows the concept space trajectory of 10, optimizing Not Male, which we intuitively expect to have a stronger concept signal, although it is intractable to compute since the DGP is unknown. Similar to our synthetic experiments, we see a transition in the concept space trajectory of the compositional class corresponding to the model disentangling the concepts. After this transition, concept learning begins but we see biases, where we see a trade-off of moving in the right direction towards one concept degrades the other one as seen in Fig. 4 (b). The visual generations in Fig. 14 confirm that our findings are not merely a result of an ill-calibrated classifier model. However, in this case, we do not observe full out-of-distribution (OOD) generalization at 1M gradient steps. We expect this task of generating (NOT Male, Smiling) is inherently harder than our synthetic setup. We note that the goal of this experiment is not to show good compositional generalization but it is more on confirming that our qualitative findings generalize to real data without touching the model or training method.
@@
C.4 Experiments with 3D concept space (Fig. 15)
@@
To further verify our findings in Sec. 4, we explore a setup where the concept conditioning specifies three concepts: color, size and background color. Fig. 15 illustrates two example scenarios in which the concept signal in color and background color are varied, respectively resulting in cases of success and failure of OOD generalization. The length of edges of the cuboids represent their concept signal magnitude. In Fig. 15 (a), we see that the object color has a strong concept signal and this is reflected in the concept space trajectories as this direction being split first. In this case we see, similar to the blue generalization curve in Fig. 4 (b) and Fig. 13 (rightmost panel), a slow generalization process for the compositional class 111. Similarly to our findings in Sec. 4.2, we observe that the 011 class initially undergoes concept memorization for class 010, which shares the two stronger concept signals color and background color, and shows a transition where it leaves this phase. In Fig. 15 (b), we see a case where out-of-distribution (OOD) generalization did not succeed within the given 80K gradient steps. In this case, we see two classes, 011 and 111, which are not correctly learned, and thus suspect that the model did not learn to generalize the background color. However, the concept space trajectories suggest that the capability is present.
Figure 15: Concept Learning Geometry underlies emergence Success and Failure modes of out-of-distribution (OOD) generalization. (a) A success case where color carries the strongest concept signal. We see aspects of concept memorization from the class 011. (b) A failure case where the background color didn’t generalize. In this case, the model does not produce the right background color while the concept space suggest that this capability is present.
C.5 Experiments with classifier free guidance
@@
An implication of our conclusions in Sec. 4.4 is that before the model has passed the transition point where the model has the capability to compose concepts, the model should not be able to generate small blue circles no matter how well “prompted”. Here, instead of prompting, we explore Classifier Free Guidance (CFG) [95] to see if our findings apply to a conditional diffusion model trained with CFG. In Fig. 16, we see that even the models with CFG show this transition from concept memorization to out-of-distribution (OOD) generalization. In a scenario where there isn’t a sharp acquisition of the capability to compositionally generalize, we would expect the sharp transition to disappear with CFG scale.
@@
Figure 16: Learning Dynamics in Concept space for Classifier-free Guidance We show the learning dynamics in the concept space for all the classes (00, 01, 10, 11). The generalization task (rightmost) shows a sharp transition from concept memorization to out-of-distribution (OOD) generalization independently of the Classifier-Free Guidance (CFG) scale.
C.6 Patching the embedding module
@@
Assuming a model does not have the capability to generate images from an OOD concept class, we simply swap the embedding module for processing the conditioning vector $h$ with that of the last checkpoint. One possible interpretation for this method is that the embedding module disentangles the concepts, i.e., generates a representation for each concept, while the U-Net [93] then utilizes such representations. This would imply that the U-Net [93] already learns how to utilize concept representations early during training, while further gradient steps lead to more robust concept representations.
@@
We test our approach on 5 random initialization seeds. Results are shown in Fig. 17 and we find that for some seeds, we are able to elicit the target behavior at around the same time in which overprompting and linear interventions also elicit the target behavior; for other seeds, this is not the case. Interestingly, we find that our method works well for seeds with “stable” learning dynamics, in that once the model reaches $\sim 100\%$ accuracy, it converges and stays at such accuracy, while some seeds have “unstable” learning dynamics in that their accuracies oscillate.

### v2

@@ B.1 Synthetic data
Fig. 13 illustrates two datasets with varying concept signals. Fig. 13 (left) illustrates a case where the color concept signal is stronger than the size concept signal, while Fig. 13 (right) illustrates the inverse case where the size concept signal is stronger.
B.2 Model Details
@@
Classifier Free Guidance At training time we drop the conditioning to a $-1.0$ filled null vector $\phi$ with probability 0.2. This unconventional choice was made instead of the dropping it to the null vector since the null vector was near the interpolation limit of our model in the color subspace of the conditioning. The $w_{cfg}$ parameter is used to estimate the noise in the diffusion process by: $\hat{\epsilon}_{t-1}=f_{\theta}(x_{t}|\phi)+w_{cfg}*(f_{\theta}(x_{t}|h)-f_{\theta}(x_{t}|\phi))$, where $\theta$ denotes parameters of the network $f$.
@@
Diffusion Process Hyperparameters We used the continuous time variational diffusion model [90] in order to (i) keep the sampling step parameter T an inference time hyperparameter and (ii) to allow the model to the adjust its optimal noise schedule for these images, especially since our synthetic images are expected to be different in SNR from natural images. In particular, we initialize our network with $\gamma_{i}=-5.0$ and $\gamma_{f}=10.0$ (see [90] for the definition of $\gamma$) and use a learned linear schedule of $\gamma(t)$. We use a reconstruction loss corresponding to a negative log likelihood from a standard normal with $\sigma=10^{-3}$ centered at the data for the first step of the diffusion process.
@@
Evaluation probe Details. We used a U-Net [95] backbone with 64 output channels followed by a max pooling layer and $n\times$ 1-layer MLP classifier for each of the $n$ classes to estimate each concept of an image independently. We sample from the same data distribution but with maximal data diversity, i.e., with $s_{i}$ values in App. B.1 maximized within the range allowing perfect classification (no overlap of $z$ between classes). We sample 4096 images per class from the DGP and train the classifier for 10K gradient steps with AdamW [122] and achieve a 100% accuracy on the held out test set. At evaluation phase, we average the classifier softmax output over 5 data generation / model initialization seeds and 32 inference samples to construct the concept space representation of the generations.
@@
Hyperparameter Search We conducted a hyperparameter search, testing batch sizes from 32 to 256, number of channels per layer from 64 to 512, learning rates between $10^{-4}$ and $10^{-3}$, the number of steps in the diffusion process from 100 to 400, weight decay between $3\times 10^{-3}$ and $5\times 10^{-2}$, and model weight initialization scale between $\mathcal{N}(0,0.003)$ and $\mathcal{N}(0,1)$. We also tried the Adam optimizer [123] with $\beta_{1}=0.9$, $\beta_{2}=0.99$, and weight decay of $10^{-5}$. No qualitative change were found in our results.
Computational Details. We implement our models in PyTorch [124]. The diffusion model was trained on four Nvidia A100 GPUs and 64 CPUs for the data generating process. A standard model run (e.g., in Sec. 4.2) took $\sim$ 20 minutes on a single NVIDIA A100 40GB GPU. The CelebA runs took $\sim$ 24 hours on the same GPU. The full research required roughly 500 GPU hours, while the experiments in the paper would require roughly 100 GPU hours.
B.3 Code Availability
Appendix C Details on Alternative Protocols for Eliciting Model Capabilities
@@
Our model is trained on a distribution of conditioning vectors centered around a class dependent mean: $p^{i}(z_{j})=\mathcal{U}(m^{i}_{j}-s_{j},m^{i}_{j}+s_{j})$ for each class $i$. We prompt the model with conditioning vectors extrapolated in the direction $\vec{m_{1}}-\vec{m_{0}}$. For instance, assuming the red conditioning was $(0.6,0.4,0.4)$ and the blue conditioning was $(0.4,0.4,0.6)$, we “prompt” the model with $(0.2,0.2,0.8)$ for blue. In practice, we use 5 conditionings $(0.4,0.4,0.6)$, $(0.35,0.35,0.65)$, $(0.25,0.25,0.75)$, $(0.15,0.15,0.85)$, $(0.05,0.05,0.95)$, and report the maximum joint accuracy.
@@
Appendix D Additional Results
D.1 Loss and Accuracy versus Concept Space (Fig. 14)
@@ Appendix D Additional Results
In Fig. 14, we plot loss, accuracy, and OOD generalization trajectory in concept space of the models for a training run in Fig. 6. The time at which the model acquires a capability to manipulate the concept color is not evident from the loss or accuracy curves; however, in concept space, one can identify when the trajectory deviates from concept memorization and thus when the compositional ability has emerged. Benchmarking generative models is a challenging task, still often involving humans in the loop [105, 106]. Our concept space framework suggests that benchmarking out-of-distribution (OOD) generalization can potentially be reduced to monitoring the learning trajectory in concept space.
@@
Figure 14: Loss vs. accuracy vs. concept space. (left) The training and validation loss on ID classes during training. (center) The accuracy on the training classes and the test (OOD generalization) class. (right) Concept space trajectory of the generalization class. Loss and accuracy do not always intuitively reflect what capabilities the model has acquired during training. However, as one can see in the rightmost panel, the point of sudden turn in concept space corresponds to when the capability has emerged, i.e., the moment when well defined prompting protocols can elicit the desired output from the model, which we indicate using the pink star (50% capability).
D.2 Additional trajectories of learning dynamics (Fig. 15)
@@
Here we show all concept space trajectories for the experiments mentioned in Fig. 4 (a,b), for all classes and color concept signal levels. We find asymmetric behavior for the 01 class and the 10 class when adjusting the color concept signal level. The dynamics of the generations in the training set matches our intuitions of concept signal as discussed in the main text. At low color concept signal, we observe that the dynamics fit size first for both 00 and 10, and afterwards find their correct colors.
@@
Figure 15: Concept space dynamics for all classes (00, 01, 10, 11). The experiment is identical as in Fig. 4. The [0,1) normalized color concept signal is color coded in every trajectory. Two training data trajectories are shown in gray in the last panel to illustrate concept memorization.
D.3 Error Quantification (Fig. 16, Fig. 17)
We show the standard error of the mean (s.e.m.) of concept space trajectories of Fig. 4 and Fig. 15 in Fig. 16 (a). Fig. 16 (b) illustrates the $50\%$ capability point across 4 different model initialization seeds. Fig. 16 (c) illustrates the $50\%$ capability point across different color concept signal level. Overall, our results are consistent over random seeds and concept signal levels.
Figure 16: Error quantification on concept space trajectories. (a) Concept space trajectories and their standard error of the mean (n=5) for different concept signal levels. (b) Concept space trajectories for 4 different seeds. The pink stars denote the point of $50\%$ capability. (c) Concept space trajectories for different color concept signal level. The stars again denote the $50\%$ capability point and the $25\%-75\%$ capability regions are in solid lines.
We show the standard deviation of compositional generalization accuracy for different prompting methods in Fig. 17. We find that the standard deviation is large for the naive prompting method, while linear latent interpolation and overprompting extract the hidden capability robustly across seeds, resulting in a small standard deviation.
Figure 17: Standard deviation of compositional generalization accuracy. We show the compositional generalization accuracy and shade the $1\sigma$ region for (a) Naive prompting (b) Linear latent intervention (c) Overprompting
D.4 Additional Experiments with real data, CelebA (Fig. 18)
To assess our findings on a more realistic dataset, we ran experiments on the CelebA [98] dataset. The experiments in Fig. 7 and Fig. 18, share the same experiment hyperparameters, described below.
@@
We selected the Female and Smiling features as the two concept dimensions to explore. This choice was motivated by the relatively balanced number of samples in all four classes (00=(Male, Not Smiling), 01=(Male, Smiling), 10=(Female, Not Smiling), 11=(Female, Smiling)), from which we randomly sampled 30K images in each class. To construct the concept space, we trained a fully convolutional network with a average pooling layer followed by a classification MLP head to classify the two attributes with independent cross entropy losses. (See B.2). We trained the classifier with the AdamW [122] optimizer with learning rate $10^{-3}$ and weight decay $10^{-5}$ for 10K gradient steps. The classifier achieved a final accuracy of respectively $95\%$ and $97\%$ on the held out validation set, which was $10\%$ of the entire dataset.
@@
We trained the same diffusion model (See App. B.2) from our synthetic experiments on 64x64 resized images from the classes (00, 01, 10) and assessed the out-of-distribution (OOD) generalization to 11. We used a color jitter of 0.1 for brightness, contrast and saturation and randomly flipped the images horizontally. As the class attributes are categorical, they were one hot encoded and concatenated to an input conditioning vector of 4 dimensions. The diffusion model was trained for $2\times 10^{6}$ gradient steps with a batch size of 64 with the same optimizer as in the main experiments.
@@
The concept space dynamics of generations from the in distribution conditioning and OOD conditioning are shown in Fig. 18. We find similar observations from the concept space trajectories as in our synthetic experiments. Initially, the images corresponding to the class 11 follows the concept space trajectory of 10, optimizing Female, which we intuitively expect to have a stronger concept signal, although it is intractable to compute since the DGP is unknown. Similar to our synthetic experiments, we see a transition in the concept space trajectory of the compositional class corresponding to the model disentangling the concepts. After this transition, concept learning begins but we observe a substantial bias towards the training distribution. We see a trade-off of moving in the right direction towards one concept degrades the other one similarly to the observations in Fig. 4 (b) for high color concept signal levels. The visual generations in Fig. 18 confirm that our findings are not merely a result of an ill-calibrated classifier model. However, in this case, we do not observe full out-of-distribution (OOD) generalization at 2M gradient steps. We expect this task of generating (Female, Smiling) is inherently harder than our synthetic setup. We note that the goal of this experiment is not to show good compositional generalization but it is more on confirming that our qualitative findings generalize to real data without touching the model or training method.
@@
D.5 Experiments with 3D concept space (Fig. 19, Fig. 20)
@@
To further verify our findings in Sec. 4, we explore a setup where the concept conditioning specifies three concepts: color, size and background color. Fig. 19 illustrates two example scenarios in which the concept signal in color and background color are varied, respectively resulting in cases of success and failure of OOD generalization by naive prompting. The length of edges of the cuboids represent their concept signal magnitude. In Fig. 19 (a), we see that the object color has a strong concept signal and this is reflected in the concept space trajectories as this direction being split first. In this case we see, similarly to the blue generalization curve in Fig. 4 (b) and Fig. 15 (rightmost panel), a slow generalization process for the compositional class 111. Again similarly to our findings in Sec. 4.2, we observe that the 011 class initially undergoes concept memorization for class 010, which shares the two stronger concept signals color and background color, and shows a transition where it suddenly curves out from this trajectory. In Fig. 19 (b), we see a case where out-of-distribution (OOD) generalization did not succeed within the given 80K gradient steps. In this case, we see two classes, 011 and 111, which didn’t reach the right compositional class via naive prompting. However, the concept space trajectories suggest that the capability to compositionally generalize background color has already emerged as visible from the sharp turns.
Figure 19: Generalization dynamics in a 3 concept setting. (a) Generalization dynamics when color carries the strongest concept signal. (b) Generalization dynamics when size carries the strongest concept signal. We observe sharp turns corresponding to learning background color, however, naive prompting cannot fully elicit compositonal generalization.
Fig. 20 quantifies the accuracy of each concept separately. Fig. 20 (a) shows that stronger concept signal accelerates the corresponding concept learning. Fig. 20 (b) shows that other concepts’ learning speed can vary, albeit smoothly, when changing a single concept signal while the corresponding concept is the most affected.
Figure 20: Accuracy on individual concepts in a 3 concept setting. a) We show the dynamics of color and background color accuracy during model training, depending on different concept signal levels b) We show the average accuracy until the end of training as a metric of learning speed for different concepts, as we change a single concept signal level.
D.6 Experiments with classifier free guidance (Fig. 21)
@@
An implication of our conclusions in Sec. 4.4 is that before the model has passed the transition point where the model has the capability to compose concepts, the model should not be able to generate small blue circles no matter how well “prompted”. Here, instead of prompting, we explore Classifier Free Guidance (CFG) [97] to see if our findings apply to a conditional diffusion model trained with CFG. In Fig. 21, we see that even the models with CFG show this transition from concept memorization to out-of-distribution (OOD) generalization. In a scenario where there isn’t a sharp acquisition of the capability to compositionally generalize, we would expect the sharp transition to disappear with CFG scale. However, as we observe in Fig. 21, the sharp turn in concept space still exists with CFG, perhaps even more pronounced than baseline experiments. These results suggest that the acquisition of the compositional generalization ability is required for CFG to enhance OOD generalization.
@@
Figure 21: Learning dynamics in Concept space for classifier-free Guidance We show the learning dynamics in the concept space for all the classes (00, 01, 10, 11). The generalization task (rightmost) shows a sharp transition from concept memorization to OOD generalization independently of the classifier-free guidance scale.
D.7 Patching the embedding module (Fig. 22)
@@
We test an additional elicitation method where we swap the conditioning vector ($h$)’s embedding module with that of the last checkpoint. An interpretation of this method is that the embedding module disentangles the concepts, i.e., generates a representation for each concept, while the U-Net [95] then utilizes such representations. This would imply that the U-Net [95] learns how to utilize concept representations early during training, while many more gradient steps are needed for the concepts to be disentangled when naively prompted.
@@
We test our approach on 5 random initialization seeds. Results are shown in Fig. 22. We find that for some seeds, we are able to elicit the target behavior at around the same time in which overprompting and linear interventions also elicit the target behavior; for other seeds, this is not the case. These results demonstrates that the role of the embedding module and the U-Net might be separated during training for some runs without an explicit term enhancing this separation of roles.